OpenStudy (anonymous):

Gemoetry Question If ST (14x) = TU (5x + 32) .. what would be the value of X? Then the value of ST? Please and thank you...Ohh and can you please explain?

OpenStudy (anonymous):

since 14x=5x+32 solve for x x=32/9 the value of ST = 14(32/9)=49.777...

OpenStudy (anonymous):

Seems good to me, steps attached.